OpenStudy (smarty):

What are the rules for dividing,multiplying,subtracting, and adding with exponents?

OpenStudy (anonymous):

http://www.sglearnonline.com/exponents.html

OpenStudy (anonymous):

x^3 times x^4 + x^7. When multiplying, you add the exponents. x^4/x^3 + x. You subtract. And you can't add or subtract x^3 and x^4. You can only add and subtract when they have the same exponent. Ex) x^3 + x^3 = 2x^3. Does this make sense?

OpenStudy (anonymous):

*****x^3 times x^4 EQUALS x^7. and x^4/x^3 EQUALS x.

OpenStudy (anonymous):

multiplying:\[x^m\times~ x^n=x^{m+n}\]dividing:\[\frac{x^m}{x^n}=x^{m-n}\]

OpenStudy (smarty):

what if it is 3x^2 X 2x^4 Will it be 6x^6

OpenStudy (anonymous):

\[3x^2\times2x^4~~~\rightarrow~~~3(x^2)\times2(x^4)~~~\rightarrow~~~6(x^{2+4})~~~\rightarrow~~~6x^6\]

OpenStudy (smarty):

Also addition and subtraction examples? @yummydum

OpenStudy (anonymous):

yes!

OpenStudy (smarty):

Also what if it is: 5E^11/6E^3 How do i solve that? I know the subtraction, but what about the whole numbers?

OpenStudy (anonymous):

the whole numbers stay as they are if they cant be simplified...so in this case your answer would be:\[5E^8\over6E\]

OpenStudy (smarty):

hmk

OpenStudy (anonymous):

when subtracting the exponents in a division problem, the resulting number, if positive is put on the variable in the numerator, and if negative, is put on the variable in the denominator

OpenStudy (anonymous):

the resulting exponent*

OpenStudy (smarty):

what about 1.01E^2+2.00E^7

OpenStudy (smarty):

@yummydum

OpenStudy (anonymous):

mm.....i think @hero should do this